I cannot figure out how to set up this problem... I created a chart with the information, but just don't feel like I have enough to solve! Can you help?
Airplane A travels 2800 km at a certain speed. Airplane B travels 2000 km at a speed that is 50 km/h faster than Plane A in 3 hours less time. Find the speed of each.

Answer by mananth(16946)

Airplane A travels 2800 km at a certain speed. Airplane B travels 2000 km at a speed that is 50 km/h faster than Plane A in 3 hours less time. Find the speed of each.
...
Airplane A travels 2800 km at x mph
Airplane B travels 2000 km at x+50 mph
...
Time plane A - time plane B = 3 hours
...
t=d/r
..
2800/x - 2000/(x+50)=3
LCD = x(x+50)
(2800(x+50)-2000x)/x(x+50)=3
2800x+140000=2000x = 3x(x+50)
800x+140000=3x^2+150x
3x^2-650x-140000=0
3x^2-1050x+400x-140000=0
3x(x-350)+400(x-350)=0
(x-350)(3x+400)=0
x = 350 mph the speed of plane A
speed of plane B = 350 +50 = 400 mph
...
m.ananth@hotmail.ca